# Dusit Niyato, , Bo Ai

### v1

@@ Appendix A Useful Results
$\mathbb{E}\{\mathbf{v}_{k}^{H}\hat{\mathbf{g}}_{l}\hat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})+p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})+p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$ (41)
@@ Appendix B Proof of Theorem 1
$\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}=\mathbb{E}\{\hat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}\overset{(a)}{=}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})$ (37)
=$\overset{(b)}{=}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{R}}_{k}),$
@@
where step (a) follows from (32) and step (b) is based on the independent feature between the $\tilde{\mathbf{g}}_{k}$ and $\hat{\mathbf{g}}_{k}$ and zero-mean feature of $\tilde{\mathbf{g}}_{k}$. Besides, we define $\bar{\mathbf{R}}_{k}\triangleq\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\}=\bar{\mathbf{G}}_{kk}+p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}$ where $\bar{\mathbf{G}}_{kk}=\mathbb{E}\{\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\}=\mathrm{diag}[\bar{\mathbf{G}}_{1kk},\dots,\bar{\mathbf{G}}_{Mkk}]$ due to the independent characteristics of $\{\varphi_{mk}\}$ for different AP-UE pairs, that is $\mathbb{E}\{e^{j\theta_{mk}}e^{-j\theta_{nl}}\}=\mathbb{E}\{e^{j\theta_{mk}}\}\mathbb{E}\{e^{-j\theta_{nl}}\}=0$ if $m\neq n$ or $k\neq l$. Meanwhile, it can be easily verified that $\mathbb{E}\{e^{j\theta_{mk}}\}=0$ according to the random distribution of $\theta_{mk}\sim\mathcal{U}[-\pi,\pi]$.
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}|\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})|^{2}}{\sum_{l=1}^{K}{p_{l}\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})\}}-p_{k}|\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})|^{2}+\sigma^{2}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\})}$ (44)
@@
$\overline{\widetilde{\mathrm{SINR}}}_{k}^{\mathrm{LSFD}}=\frac{p_{k}|\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{R}}_{mk})}|^{2}}{\left(\begin{array}[]{c}\sum_{l=1}^{K}{\sum_{m=1}^{M}{p_{l}\varphi_{mkl}}}+\sum_{l\in\mathcal{P}_{k}}{\sum_{m=1}^{M}{p_{l}\upsilon_{mkl}}}+\sum_{l\in\mathcal{P}_{k}}{p_{l}(|\sum_{m=1}^{M}{\gamma_{mkl}}|^{2}-\sum_{m=1}^{M}{|\gamma_{mkl}|^{2}})}\\ -p_{k}|\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{R}}_{mk})}|^{2}+\sigma^{2}\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}\bar{\mathbf{R}}_{mk}\mathbf{W}_{mk}^{H})}\\ \end{array}\right)}$ (40)
As for $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$, if $l\notin\mathcal{P}_{k}$, $\hat{\mathbf{g}}_{k}$ and $\mathbf{g}_{l}$ are independent. We denote $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$ as
[…]
@@
where $\hat{\mathbf{g}}_{k,\mathrm{NLoS}}=\sqrt{p_{k}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}(\mathbf{y}_{k}^{p}-\bar{\mathbf{y}}_{k}^{p})$. Then, we can compute terms in (41) in closed-form by expanding them as $\mathbb{E}\left\{aceg\right\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$, $\mathbb{E}\{adfg\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$, $\mathbb{E}\{bceh\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\hat{\mathbf{R}}_{k})$, and $\mathbb{E}\{bdfh\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\check{\mathbf{g}}_{l}\check{\mathbf{g}}_{l}^{H}\}\mathbf{W}_{k}\mathbb{E}\{\hat{\mathbf{g}}_{k,\mathrm{NLoS}}\hat{\mathbf{g}}_{k,\mathrm{NLoS}}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\hat{\mathbf{R}}_{k})$. The other terms are all zero, which can be easily verified. In summary, for $l\notin\mathcal{P}_{k}$, we have $\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{v}_{k}\}=\mu_{kl}$. If $l\in\mathcal{P}_{k}\backslash\{k\}$, $\hat{\mathbf{g}}_{k}$ and $\mathbf{g}_{l}$ are not independent since $\mathbf{y}_{k}^{p}$ contains $\mathbf{g}_{l}$. By following the method in [29, Appendix D], we can construct $\hat{\mathbf{g}}_{k}$ and $\hat{\mathbf{g}}_{l}$ as $\hat{\mathbf{g}}_{k}=\bar{\mathbf{g}}_{k}+\sqrt{p_{k}\tau_{p}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-\frac{1}{2}}\mathbf{q}$ and $\hat{\mathbf{g}}_{l}=\bar{\mathbf{g}}_{l}+\sqrt{p_{l}\tau_{p}}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-\frac{1}{2}}\mathbf{q}$, respectively, where $\mathbf{q}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\mathbf{I}_{MN})$. We formulate $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$ as
$\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}=\mathbb{E}\{\mathbf{v}_{k}^{H}\hat{\mathbf{g}}_{l}\hat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}+\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}.$ (42)
For $\mathbb{E}\{\mathbf{v}_{k}^{H}\hat{\mathbf{g}}_{l}\hat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$, we have
@@
By expanding terms in (43), we can compute them in closed-form as $\mathbb{E}\{aceg\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$, $\mathbb{E}\{adfg\}=p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$, $\mathbb{E}\left\{bceh\right\}=p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$, and $\mathbb{E}\{bdfh\}\overset{(a)}{=}p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\hat{\mathbf{R}}_{k})+p_{k}p_{l}\tau_{p}^{2}|\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})|^{2}$ where step (a) is according to [29, Eq. 80] by letting $(\mathbf{R}_{x}^{H})^{\frac{1}{2}}=\mathbf{\Psi}_{k}^{-\frac{1}{2}}\check{\mathbf{R}}_{k}$, $\mathbf{R}_{y}^{{\frac{1}{2}}}=\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-\frac{1}{2}}$, and, $\mathbf{B}=\mathbf{W}_{k}^{H}$, respectively. The other terms are zero due to the circular symmetry characteristic of $\mathbf{q}$ and the independent characteristic of $\{\theta_{mk}\}$ for different AP-UE pairs. In summary, for $l\in\mathcal{P}_{k}\backslash\{k\}$, we can compute $\mathbb{E}\{\mathbf{v}_{k}^{H}\hat{\mathbf{g}}_{l}\hat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$ in closed-form as in (41) on the top of the next page. As for $\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$, relying on the independence feature between $\hat{\mathbf{g}}_{k}$ and $\tilde{\mathbf{g}}_{l}$, we can compute it in closed form
$\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}=\mathbb{E}\{|(\mathbf{W}_{k}\hat{\mathbf{g}}_{k})^{H}\tilde{\mathbf{g}}_{l}|^{2}\}$ (42)
$=p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{C}_{l}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})+\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{C}_{l}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$
@@
by applying [29, Lemma 4] with $\mathbf{R}_{x}=p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}$, $\mathbf{R}_{y}=\mathbf{C}_{l}$, and $\mathbf{B}=\mathbf{W}_{k}^{H}$, respectively. By plugging (41) and (42) into (42), we can compute $\mathbb{E}\left\{\mathbf{v}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{v}_{k}\right\}$ in closed-form as $\mu_{kl}+\varepsilon_{kl}$ by applying the fact that $p_{l}\tau_{p}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}=\check{\mathbf{R}}_{l}-\mathbf{C}_{l}$. For $l=k$, besides terms $\mu_{kl}+\varepsilon_{kl}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$, there are also additional terms $\mathbb{E}\{acfh\}=p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{kk})\mathrm{tr(}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$ and $\mathbb{E}\{bdeg\}=\!p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})\mathrm{tr(}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$. Moreover, the computation of $\mathbb{E}\{aceg\}=\mathbb{E}\{\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\bar{\mathbf{g}}_{k}\}=\mathbb{E}\{|\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}|^{2}\}$ is quite tricky since there are $M$ different phase-shifts in $\bar{\mathbf{g}}_{k}$ from $M$ APs. By constructing $\mathbf{W}_{k}$ into the block matrix form, we can decompose $\mathbb{E}\{|\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}|^{2}\}$ as
$\mathbb{E}\{|\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}|^{2}\}=\sum_{\mathbf{m}}{\mathbb{E}\{\bar{\mathbf{g}}_{m_{1}k}^{H}e^{-j\theta_{m_{1}}k}\mathbf{W}_{k,m_{2}m_{1}}^{H}\bar{\mathbf{g}}_{m_{2}k}e^{j\theta_{m_{2}}k}}$ (43)
$\bar{\mathbf{g}}_{m_{3}k}^{H}e^{-j\theta_{m_{3}}k}\mathbf{W}_{k,m_{3}m_{4}}\bar{\mathbf{g}}_{m_{4}k}e^{j\theta_{m_{4}}k}\},$

### v2

@@ Appendix A Useful Results
[…]
$\mathbb{E}\{\mathbf{v}_{k}^{H}\widehat{\mathbf{g}}_{l}\widehat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})+p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})+p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$ (41)
@@ Appendix B Proof of Theorem 1
$\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{k}\}=\mathbb{E}\{\widehat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{g}_{k}\}\overset{(a)}{=}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})$ (37)
=$\overset{(b)}{=}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{R}}_{k}),$
@@
where step (a) follows from (32) and step (b) is based on the independent feature between the $\tilde{\mathbf{g}}_{k}$ and $\widehat{\mathbf{g}}_{k}$ and zero-mean feature of $\tilde{\mathbf{g}}_{k}$. Besides, we define $\overline{\mathbf{R}}_{k}\triangleq\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\}=\overline{\mathbf{G}}_{kk}+p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}$, where $\overline{\mathbf{G}}_{kk}=\mathbb{E}\{\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\}=\mathrm{diag}[\overline{\mathbf{G}}_{1kk},\dots,\overline{\mathbf{G}}_{Mkk}]$ due to the independent characteristics of $\{\varphi_{mk}\}$ for different AP-UE pairs, that is $\mathbb{E}\{e^{j\theta_{mk}}e^{-j\theta_{nl}}\}=\mathbb{E}\{e^{j\theta_{mk}}\}\mathbb{E}\{e^{-j\theta_{nl}}\}=0$ if $m\neq n$ or $k\neq l$. Meanwhile, it can be easily verified that $\mathbb{E}\{e^{j\theta_{mk}}\}=0$ according to the random distribution of $\theta_{mk}\sim\mathcal{U}[-\pi,\pi]$.
@@
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}|\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})|^{2}}{\sum_{l=1}^{K}{p_{l}\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})\}}-p_{k}|\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})|^{2}+\sigma^{2}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\})}$ (44)
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}|\mathbf{w}_{k}^{H}\mathrm{vec(}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})|^{2}}{\mathbf{w}_{k}^{H}\left[\sum_{l=1}^{K}{p_{l}\mathbb{E}\{[(\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})]\}-p_{k}|\mathbf{w}_{k}^{H}\mathrm{vec(}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})|^{2}+(\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})}\right]\mathbf{w}_{k}}$ (45)
@@
$\overline{\widetilde{\mathrm{SINR}}}_{k}^{\mathrm{LSFD}}=\frac{p_{k}|\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{R}}_{mk})}|^{2}}{\left(\begin{array}[]{c}\sum_{l=1}^{K}{\sum_{m=1}^{M}{p_{l}\varphi_{mkl}}}+\sum_{l\in\mathcal{P}_{k}}{\sum_{m=1}^{M}{p_{l}\upsilon_{mkl}}}+\sum_{l\in\mathcal{P}_{k}}{p_{l}(|\sum_{m=1}^{M}{\gamma_{mkl}}|^{2}-\sum_{m=1}^{M}{|\gamma_{mkl}|^{2}})}\\ -p_{k}|\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{R}}_{mk})}|^{2}+\sigma^{2}\sum_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}\overline{\mathbf{R}}_{mk}\mathbf{W}_{mk}^{H})}\\ \end{array}\right)}$ (40)
@@
As for $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$, if $l\notin\mathcal{P}_{k}$, $\widehat{\mathbf{g}}_{k}$ and $\mathbf{g}_{l}$ are independent. We denote $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$ in (38) on the top of the next page, where $\widehat{\mathbf{g}}_{k,\mathrm{NLoS}}=\sqrt{p_{k}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}(\mathbf{y}_{k}^{p}-\overline{\mathbf{y}}_{k}^{p})$. Then, we can compute terms in (38) in closed-form by expanding them as $\mathbb{E}\left\{aceg\right\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$, $\mathbb{E}\{adfg\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$, $\mathbb{E}\{bceh\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\widehat{\mathbf{R}}_{k})$, and $\mathbb{E}\{bdfh\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\check{\mathbf{g}}_{l}\check{\mathbf{g}}_{l}^{H}\}\mathbf{W}_{k}\mathbb{E}\{\widehat{\mathbf{g}}_{k,\mathrm{NLoS}}\widehat{\mathbf{g}}_{k,\mathrm{NLoS}}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\widehat{\mathbf{R}}_{k})$. The other terms are all zero, which can be easily verified. In summary, for $l\notin\mathcal{P}_{k}$, we have $\mathbb{E}\{\mathbf{v}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{v}_{k}\}=\mu_{kl}$. If $l\in\mathcal{P}_{k}\backslash\{k\}$, $\widehat{\mathbf{g}}_{k}$ and $\mathbf{g}_{l}$ are not independent since $\mathbf{y}_{k}^{p}$ contains $\mathbf{g}_{l}$. By following the method in [31, Appendix D], we can construct $\widehat{\mathbf{g}}_{k}$ and $\widehat{\mathbf{g}}_{l}$ as $\widehat{\mathbf{g}}_{k}=\overline{\mathbf{g}}_{k}+\sqrt{p_{k}\tau_{p}}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-\frac{1}{2}}\mathbf{q}$ and $\widehat{\mathbf{g}}_{l}=\overline{\mathbf{g}}_{l}+\sqrt{p_{l}\tau_{p}}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-\frac{1}{2}}\mathbf{q}$, respectively, where $\mathbf{q}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\mathbf{I}_{MN})$. We formulate $\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}$ as
$\mathbb{E}\{|\mathbf{v}_{k}^{H}\mathbf{g}_{l}|^{2}\}=\mathbb{E}\{\mathbf{v}_{k}^{H}\widehat{\mathbf{g}}_{l}\widehat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}+\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}.$ (39)
For $\mathbb{E}\{\mathbf{v}_{k}^{H}\widehat{\mathbf{g}}_{l}\widehat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$, we have
@@
By expanding terms in (40), we can compute them in closed-form as $\mathbb{E}\{aceg\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$, $\mathbb{E}\{adfg\}=p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$, $\mathbb{E}\left\{bceh\right\}=p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$, and $\mathbb{E}\{bdfh\}\overset{(a)}{=}p_{l}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\widehat{\mathbf{R}}_{k})+p_{k}p_{l}\tau_{p}^{2}|\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})|^{2}$ where step (a) is according to [31, Eq. 80] by letting $(\mathbf{R}_{x}^{H})^{\frac{1}{2}}=\mathbf{\Psi}_{k}^{-\frac{1}{2}}\check{\mathbf{R}}_{k}$, $\mathbf{R}_{y}^{{\frac{1}{2}}}=\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-\frac{1}{2}}$, and $\mathbf{B}=\mathbf{W}_{k}^{H}$, respectively. The other terms are zero due to the circular symmetry characteristic of $\mathbf{q}$ and the independent characteristic of $\{\theta_{mk}\}$ for different AP-UE pairs. In summary, for $l\in\mathcal{P}_{k}\backslash\{k\}$, we can compute $\mathbb{E}\{\mathbf{v}_{k}^{H}\widehat{\mathbf{g}}_{l}\widehat{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$ in closed-form in (41) on the top of the next page. As for $\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}$, relying on the independence feature between $\widehat{\mathbf{g}}_{k}$ and $\tilde{\mathbf{g}}_{l}$, we can compute it in closed form
$\mathbb{E}\{\mathbf{v}_{k}^{H}\tilde{\mathbf{g}}_{l}\tilde{\mathbf{g}}_{l}^{H}\mathbf{v}_{k}\}=\mathbb{E}\{|(\mathbf{W}_{k}\widehat{\mathbf{g}}_{k})^{H}\tilde{\mathbf{g}}_{l}|^{2}\}$ (42)
$=p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{C}_{l}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})+\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{C}_{l}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$
@@
by applying [31, Lemma 4] with $\mathbf{R}_{x}=p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}$, $\mathbf{R}_{y}=\mathbf{C}_{l}$, and $\mathbf{B}=\mathbf{W}_{k}^{H}$, respectively. By plugging (41) and (42) into (39), we can compute $\mathbb{E}\left\{\mathbf{v}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{v}_{k}\right\}$ in closed-form as $\mu_{kl}+\varepsilon_{kl}$ by applying the fact that $p_{l}\tau_{p}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{l}=\check{\mathbf{R}}_{l}-\mathbf{C}_{l}$. For $l=k$, besides terms $\mu_{kl}+\varepsilon_{kl}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$, there are also additional terms $\mathbb{E}\{acfh\}=p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{kk})\mathrm{tr(}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})$ and $\mathbb{E}\{bdeg\}=\!p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})\mathrm{tr(}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$. Moreover, the computation of $\mathbb{E}\{aceg\}=\mathbb{E}\{\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\overline{\mathbf{g}}_{k}\}=\mathbb{E}\{|\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}|^{2}\}$ is quite tricky since there are $M$ different phase-shifts in $\overline{\mathbf{g}}_{k}$ from $M$ APs. By constructing $\mathbf{W}_{k}$ into the block matrix form, we can decompose $\mathbb{E}\{|\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}|^{2}\}$ as
$\mathbb{E}\{|\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}|^{2}\}\!\!=\!\!\sum_{\mathbf{m}}{\mathbb{E}\{\overline{\mathbf{g}}_{m_{1}k}^{H}e^{-j\theta_{m_{1}}k}\mathbf{W}_{k,m_{2}m_{1}}^{H}\overline{\mathbf{g}}_{m_{2}k}e^{j\theta_{m_{2}}k}}$ (43)
$\overline{\mathbf{g}}_{m_{3}k}^{H}e^{-j\theta_{m_{3}}k}\mathbf{W}_{k,m_{3}m_{4}}\overline{\mathbf{g}}_{m_{4}k}e^{j\theta_{m_{4}}k}\},$